# Task Preference Optimization: Improving Multimodal Large Language Models with Vision Task Alignment

### v1

@@ Abstract
Current multimodal large language models (MLLMs) struggle with fine-grained or precise understanding of visuals though they give comprehensive perception and reasoning in a spectrum of vision applications. Recent studies either develop tool-using or unify specific visual tasks into the autoregressive framework, often at the expense of overall multimodal performance. To address this issue and enhance MLLMs with visual tasks in a scalable fashion, we propose Task Preference Optimization (TPO), a novel method that utilizes differentiable task preferences derived from typical fine-grained visual tasks. TPO introduces learnable task tokens that establish connections between multiple task-specific heads and the MLLM. By leveraging rich visual labels during training, TPO significantly enhances the MLLM’s multimodal capabilities and task-specific performance. Through multi-task co-training within TPO, we observe synergistic benefits that elevate individual task performance beyond what is achievable through single-task training methodologies. Our instantiation of this approach with VideoChat and LLaVA demonstrates an overall 14.6% improvement in multimodal performance compared to baseline models. Additionally, MLLM-TPO demonstrates robust zero-shot capabilities across various tasks, performing comparably to state-of-the-art supervised models.
@@ 1 Introduction
Multimodal large language models (MLLMs) show impressive visual perception and reasoning capabilities, with applications in personal assistants [64, 71], embodied systems [20, 91], and scientific discovery [78], among others [6, 58, 13]. Considering the growing expectations of users in more accurate and detailed perception, taking the shell game as an example, further improving the generality of MLLM requires fine-grained knowledge representation beyond words, e.g. MLLM implicitly embeds how to track keys indicated by users.
Existing studies address the enhancement of perceptual granularity in MLLMs by focusing on specific visual tasks (e.g. temporal grounding, segmentation, tracking) via MLLMs. They usually fine-tune MLLMs on more task data in text format or enable MLLMs to activate the corresponding task heads. Shikra [8] applies MLLM to localization tasks, transforms object coordinates into dialogue formats, and learns them autoregressively. Meanwhile, TimeChat [72] formulates event timestamps as a text for autoregressive prediction, thereby endowing MLLMs with temporal grounding. They do improve specific task performance significantly, while at the cost of multimodal performance more or less. This is counterintuitive as seminal research proves different visual tasks are correlated and training them together often yields mutual gains [75, 50, 32]. We conjecture that the presentation of different tasks influences this training and that the conflict arises from the learning discrepancies between discrete textual tokens and visually dense predictions. Our experiments in Section 4 validate that a decoupled representation design can effectively address this issue.
To enhance the multimodal capabilities of MLLMs, we explore optimization methods to meet multiple visual task requirements in an end-to-end manner. We propose task preference optimization (TPO), incorporating visual task knowledge into the MLLMs by jointly maximizing the likelihoods of visual task estimations and multimodal dialogue. Inspired by direct preference optimization [67] (DPO) and related methods, we treat visual task annotations as human preferences in particular demands, as shown in Figure  2. DPO aids LLM (or MLLM) better in better aligning with human preferences through a binary classification that directs the model to generate responses that people prefer. Similarly, TPO enhances MLLMs’ visual sensing capabilities through differentiable task optimizations that guide MLLMs to yield dense predictions closely resembling human perception. To achieve this, TPO concretizes MLLM-specific visual perceptions into corresponding task tokens, disentangled from MLLM representation. Then it fine-tunes these task tokens and then updates the MLLM accordingly.
@@
TPO demonstrates scalability across various MLLM approaches, encompassing a wide range of visual task categories and data quantities. We validate the effectiveness of TPO within widely used MLLMs, such as LLaVA [51] and VideoChat [46], as detailed in Section 4.3. By fine-tuning these open-source MLLMs with TPO, we enhance their visual understanding capabilities and improve dialogue performance. Additionally, we explore several key spatiotemporal and spatial perception tasks, including spatial grounding, tracking, and temporal grounding. Our findings indicate that these tasks can mutually enhance each other’s performance, particularly contributing to the improvement of multimodal dialogue capabilities.
@@ 2 Related Work
Vision foundation models [96, 75, 47, 116, 83, 99, 77, 81, 88, 89, 87, 80] are designed to be adaptable for various downstream tasks through extensive pre-training on large-scale, diverse datasets. VideoPrism [116] achieves SOTA results on various video tasks by combining video-text contrastive learning and video token reconstruction using a dataset of public and proprietary videos. InternVideo2 [89] utilize masked reconstruction, cross-modal contrastive learning, and next-token prediction to enhance the model’s perceptiveness, semantic understanding, and reasoning capabilities.
@@
Based on vision foundation models, some studies [55, 120, 44, 119, 86] intend to incorporate downstream task heads into this framework and expect end-to-end task training. Unified-IO [55] builds a model that can support a diverse set of tasks across vision and language with little to no need for task-specific customizations and parameters.
Uni-Perceivers [120, 44, 119] formulate different tasks to find the maximum likelihood target for each input through the representation similarity regardless of their modality. Nevertheless, these generalist models are limited to predefined tasks and cannot support flexible, open-ended task customization based on language instructions.
@@
The effective understanding and reasoning of LMMs have attracted the attention of many researchers. Limited to its input modality, researchers expand the visual capabilities of LLM, leading to MLLMs. Seminal works, such as BLIP-2 [45], LLaVA [51], and mPLUG-Owl [101], introduce image captioning and visual question answering based on LLM by visual instruction-tuning data. Some video-based MLLMs have been proposed, such as VideoChat [46], VideoChatGPT [57], and Video-LLaMA [112], which enable LLM to gain video understanding capabilities by encoding multiple video frames and using video instruction data.
Vanilla MLLMs have achieved impressive results in visual captioning and question answering, but barely address fine-grained visual tasks, such as segmentation and temporal grounding, with precise estimations. To address this challenge, MLLMs usually take one of the three pipelines: pixel-to-sequence (P2S), and pixel-to-embedding (P2E). For P2S methods [8, 72, 90, 102, 79, 108], MLLM directly outputs the prediction results in text form. TimeChat [72] introduces a time-aware frame encoder and a sliding video Q-former in the MLLM to enhance the perception of temporal information. For P2E methods [38, 109, 4, 84, 94], MLLM compresses visual information and inputs it into the downstream decoder, which outputs the prediction results. LISA [38], NExT-Chat [109] and VideoLISA [4] introduce SAM [36] as a segmentation tool based on MLLM, using a special token as a prompt to connect MLLM and SAM. VisionLLM v2 [94] designs multiple routing tokens and super-link queries to bridge MLLM with multiple decoders.
@@
Aligning MLLM with human preferences or values is crucial for MLLM’s development. Recent works [76, 118, 106] introduce alignment approaches to MLLM, including proximal policy optimization (PPO) [73] and direct preference optimization (DPO) [67], as shown in Figure 2. They usually exploit proprietary models like GPT4-V to build visual preference datasets and then tune MLLMs using PPO or DPO, accordingly. Llava-RLHF [76] incorporates PPO into a MLLM framework Llava, argumented by image captions or question-answers (QA), while Zhou et al. [118] and Zhang et al. [114] give DPO implementations for MLLMs where visual preference data are created by GPT-4V and other open-sourced MLLMs. RLHF-V [106] collects a dense human preference in segmentation and enhances MLLMs with DPO.
@@ 3 Method
Task Preference Optimization (TPO) aims to enable a MLLM master classical visual perceptions (such as tracking, temporal grounding, etc) for better task generalization. Many multimodal reasoning tasks require precise visual cues for accurate and reliable responses. As given in Figure  3, MLLM-TPO has a typical multimodal model $M$ (consisting of a vision encoder $E$, a vision-language connector $C$, and a large language model $G$) and a task preference model $P$ with a series of visual task heads $\{H_{i}\}_{i=1...n}$ ($n$ denotes the task head number). These heads connect with MLLM using the embeddings $\{\mathbf{e}_{i}\}_{i=1...n}$ ($\mathbf{e}_{i}=G(\mathbf{v}_{i})\in\mathbb{R}^{1\times\text{C}}$) transformed from the learnable task-specific tokens $\{\mathbf{v}_{i}\}_{i=1...n}$ ($\mathbf{v}_{i}\in\mathbb{R}^{1\times\text{C}}$) via MLLM.
TPO employs a local-to-global training scheme, first adapting task heads to the MLLM and then training them jointly. Specifically, a MLLM $M$ starts to recognize $\{\mathbf{v}_{i}\}_{i=1...n}$ by updating $G$ according to user instruction, then we tune $\{\mathbf{v}_{i}\}_{i=1...n}$ and $\{H_{i}\}_{i=1...n}$ for adapting visual heads to $M$. Finally, we train both $M$ and $\{H_{i}\}_{i=1...n}$ together. During inference, MLLM-TPO can respond to users’ queries with visuals with text, and produce structured visual outputs (like masks, timestamps, trajectories, etc) when users ask (e.g. yielding time results for “find when the birthday party starts”). We detail how to form MLLM-TPO structurally and train it in the following.
@@ 3.1 Task Preference Models
The task preference model (TPM) $P$ contains a series of task tokens $\{\mathbf{v}_{i}\}_{i=1...n}$ and heads $\{H_{i}\}_{i=1...n}$. Before TPM works, the attached MLLM generates special tokens indicating task types from the input queries. Then TPM dynamically calls the task token $\mathbf{v}_{j}$ based on the special token, transforms it to the task embedding $\mathbf{e}_{j}$ via the LLM $G$ (the last hidden embedding), and feeds $\mathbf{e}_{j}$ to the corresponding visual task head $\{H_{j}\}$ for specific task predictions. Considering existing MLLMs demonstrate remarkable capabilities in common object and scene recognition, yet struggle to accurately locate things or actions, the employed task heads mainly focus on spatiotemporal localization and tracking. Specifically, we give three fundamental task types for compensating mainstream MLLMs’ gaps from expert models in visual perceptions: 1) region head, 2) temporal head, and 3) mask head. Their architectures are given blew.
  • Region Head. A two-layer multilayer perceptron (MLP) with ReLU activation is employed for the region head. It takes embeddings from LLM and regresses them to the bounding box coordinates for the spatial grounding task.
  • Temporal Head. This head is composed of a video encoder, text encoder, and temporal grounding model for moment retrieval and highlight detection. The original video and text query are input to the video and text encoder for their features respectively. Then we append the temporal task embedding after the text features. The new text features and video features output the start and end time of the corresponding moment of the query and the highlight score through the temporal grounding model.
  • Mask Head. Pixel-level tasks pose significant challenges for MLLMs due to the lack of pixel-level output capabilities in MLLMs. To this end, we introduce a specialized mask head, utilizing the image encoder, mask decoder, and memory bank components of SAM2  [69], replacing the prompt encoder with a single MLP layer called the mask adapter. Specifically, the image features extracted by the image encoder and the text features corresponding to the mask task embedding from LLM are fed into the mask decoder to produce the final mask.
Most known discriminative vision tasks can be addressed by one or a combination of the aforementioned three task heads. TPM builds the architectural foundations for leveraging existing vision annotations to enhance MLLMs.
@@ 3.2 Task Preference Optimization
TPO improves MLLMs with extra supervision from visual task heads by back-propagating gradients from heads to update MLLMs using visual task data. It enables the language model $G$ in MLLM to discriminate specific task types when users demand (task assignment). Then, TPO trains TPM $P$ via compact task representations $\mathbf{e}_{i}$ (task optimization) from $\mathbf{v}_{i}$. Lastly, we train $M$ and $P$ together, tuning $M$ for the refined spatiotemporal perception according to task preferences from $P$. Its optimization objective is given as:
@@
where $\mathbf{T}_{\text{q}}$ and $\mathbf{s}$ denote a user query (some contain specific task needs like tracking) in the form of text token sequence and the grounding truth task indicating token, respectively. $\mathbf{A}_{i}$ is the task annotation of the input visual $\mathbf{X}$, so it could be a number tuple for describing regional localization and area or mask for delineating object shape and position. Here $(\mathbf{X},\mathbf{T}_{\text{q}},\mathbf{T}_{\text{a}},\mathbf{A}_{i})$ stands for an input data tuple to MLLM-TPO. Usually, we have $(\mathbf{X},\mathbf{T}_{\text{q}},\mathbf{T}_{\text{a}})$ for typical MLLM training while $(\mathbf{X},\mathbf{T}_{\text{q}},\mathbf{A}_{i})$ for task token and head training. $\mathcal{L}_{\text{assign}}$ is the cross-entropy loss, and $\mathcal{L}_{\text{task}_{i}}$ varies according to the task and are usually regression- or classification-related losses.
To train MLLM-TPO, we propose a 3-stage local-to-global training scheme. Stage 1 learns to identify the task type based on user queries. In stage 2, we train task heads along with their corresponding task tokens, respectively. Lastly, we co-train task heads with MLLM by both task data and multimodal conversation data. Our 3-stage training strategy mitigates the risk of degrading the MLLM’s general abilities. We describe them as:
@@
We integrate task tokens and task-specific heads into the model. By training on task-specific data. The model equips the capacity with fine-grained perception and aligns between the task head and the MLLM. The region head, temporal head, mask adapter, and corresponding task tokens are trained respectively. Similarly, the LLM is updated with LoRA.

### v2

@@ Abstract
Current multimodal large language models (MLLMs) struggle with fine-grained or precise understanding of visuals although they give comprehensive perception and reasoning in a spectrum of vision applications. Recent studies either develop tool-using or unify specific visual tasks into the autoregressive framework, often at the expense of overall multimodal performance. To address this issue and enhance MLLMs with visual tasks in a scalable fashion, we propose Task Preference Optimization (TPO), a novel method that utilizes differentiable task preferences derived from typical fine-grained visual tasks. TPO introduces learnable task tokens that establish connections between multiple task-specific heads and the MLLM. By leveraging rich visual labels during training, TPO significantly enhances the MLLM’s multimodal capabilities and task-specific performance. Through multi-task co-training within TPO, we observe synergistic benefits that elevate individual task performance beyond what is achievable through single-task training methodologies. Our instantiation of this approach with VideoChat and LLaVA demonstrates an overall 14.6% improvement in multimodal performance compared to baseline models. Additionally, MLLM-TPO demonstrates robust zero-shot capabilities across various tasks, performing comparably to state-of-the-art supervised models.
@@ 1 Introduction
Multimodal large language models (MLLMs) show impressive visual perception and reasoning capabilities, with applications in personal assistants [75], embodied systems [19], and scientific discovery [7]. Considering the growing expectations of users in more accurate and detailed perception, taking the shell game as an example, further improving the generality of MLLM requires fine-grained knowledge representation beyond words, e.g. MLLM implicitly embeds how to track keys indicated by users.
Existing studies address the enhancement of perceptual granularity in MLLMs by focusing on specific visual tasks (e.g. temporal grounding, segmentation, tracking) via MLLMs. They usually fine-tune MLLMs on more task data in text format or enable MLLMs to activate the corresponding task heads. Shikra [8] applies MLLM to localization tasks, transforms object coordinates into dialogue formats, and learns them autoregressively. Meanwhile, TimeChat [70] and VideoChat-T [103] treat event timestamps as text for autoregressive prediction, thereby endowing MLLMs with temporal grounding. They do improve specific task performance significantly, while at the cost of multimodal performance more or less. This is counterintuitive as seminal research proves different visual tasks are correlated and training them together often yields mutual gains [73, 50, 31]. We conjecture that the presentation of different tasks influences this training and that the conflict arises from the learning discrepancies between discrete textual tokens and visually dense predictions. Our experiments in Section 4 validate that a decoupled representation design can effectively address this issue.
To enhance the multimodal capabilities of MLLMs, we explore optimization methods to meet multiple visual task requirements in an end-to-end manner. We propose task preference optimization (TPO), incorporating visual task knowledge into the MLLMs by jointly maximizing the likelihoods of visual task estimations and multimodal dialogue. Inspired by direct preference optimization [66] (DPO) and related methods, we treat visual task annotations as human preferences in particular demands, as shown in Figure  2. DPO aids LLM (or MLLM) in aligning with human preferences through a binary classification that directs the model to generate responses that people prefer. Similarly, TPO enhances MLLMs’ visual sensing capabilities through differentiable task optimizations that guide MLLMs to yield dense predictions closely resembling human perception. To achieve this, TPO concretizes MLLM-specific visual perceptions into corresponding task tokens, disentangled from MLLM representation. Then TPO fine-tunes these task tokens and updates the MLLM accordingly.
@@
TPO demonstrates scalability across various MLLM approaches, encompassing a wide range of visual task categories and data quantities.
We validate the effectiveness of TPO within widely used MLLMs, such as LLaVA [51] and VideoChat [46], as detailed in Section 4.3. By fine-tuning these open-source MLLMs with TPO, we enhance their visual understanding capabilities and improve dialogue performance. Additionally, we explore several key spatiotemporal and spatial perception tasks, including spatial grounding, tracking, and temporal grounding. Our findings indicate that these tasks can mutually enhance each other’s performance, particularly contributing to the improvement of multimodal dialogue capabilities.
@@ 2 Related Work
Vision foundation models [91, 73, 47, 111, 81, 94, 76, 79, 85, 78] are designed to be adaptable for various downstream tasks through extensive pre-training on large-scale and diverse datasets. VideoPrism [111] achieves the leading results in various video tasks by combining video-text contrastive learning and spatiotemporal token reconstruction using a dataset of public and proprietary videos. InternVideo2 [85] utilizes masked reconstruction, cross-modal contrastive learning, and next-token prediction to enhance the model’s perceptiveness, semantic understanding, and reasoning capabilities.
@@
Based on vision foundation models, some studies [55, 115, 44, 114, 84] intend to incorporate downstream task heads into this framework and expect end-to-end task training. Unified-IO [55] introduces a versatile model capable of handling a wide range of tasks across vision and language domains, requiring minimal or no task-specific modifications or additional parameters. Uni-Perceivers [115, 44, 114] formulate different tasks to find the maximum likelihood target for each input through the representation similarity regardless of their modality. Nevertheless, these generalist models are limited to predefined tasks and cannot support flexible open-ended task customization based on language instructions.
@@
The effective understanding and reasoning of LMMs has attracted the attention of numerous researchers. Limited to its input modality, researchers expand the visual capabilities of LLM, leading to MLLMs. Seminal works, such as BLIP-2 [45], LLaVA [51], and mPLUG-Owl [96], introduce image captioning and visual question answering based on LLM using visual instruction-tuning data. Some video-based MLLMs have been proposed, such as VideoChat [46], VideoChatGPT [57], and Video-LLaMA [107], enabling LLM to gain video understanding capabilities by encoding multiple video frames and using video instruction data.
Vanilla MLLMs have achieved impressive results in visual-related dialogues, but barely address some fine-grained tasks, such as segmentation and temporal grounding. To address this challenge, MLLMs usually take one of the following pipelines: pixel-to-sequence (P2S) and pixel-to-embedding (P2E). For P2S methods [8, 70, 86, 97, 77], MLLM directly outputs the textual predictions. TimeChat [70] introduces a time-aware frame encoder and a sliding video Q-former to enhance temporal perception. For P2E methods [37, 104, 4, 82, 89], MLLM compresses and inputs visuals into the downstream decoder for estimations. LISA [37], NExT-Chat [104], and VideoLISA [4] introduce SAM [35] as a segmentation tool, using a special token as a prompt to connect MLLM and SAM. VisionLLM v2 [89] designs routing tokens and super-link queries to bridge MLLM with multiple decoders.
@@
Aligning MLLM with human preferences or values is crucial for MLLM’s development. Recent works [74, 113, 101] introduce alignment approaches to MLLM, including proximal policy optimization (PPO) [71] and direct preference optimization (DPO) [66], as shown in Figure 2. They usually exploit proprietary models like GPT4-V to build visual preference datasets and then tune MLLMs using PPO or DPO. Llava-RLHF [74] incorporates PPO into a MLLM framework Llava, argumented by image captions or question-answers (QA), while Zhou et al. [113] and Zhang et al. [109] give DPO implementations for MLLMs where visual preference data are created by GPT-4V and other open-sourced MLLMs. RLHF-V [101] collects a dense human preference in segmentation and enhances MLLMs with DPO.
@@ 3 Method
Task Preference Optimization (TPO) aims to enable MLLMs to master classical visual perceptions (such as tracking, temporal grounding, etc.) for better task generalization, regarding that many multimodal reasoning tasks require precise visual cues for accurate and reliable responses. As shown in Figure  3, MLLM-TPO has a typical multimodal model $M$ (consisting of a vision encoder $E$, a vision-language connector $C$, and a large language model $G$) and a task preference model $P$ with a series of visual task heads $\{H_{i}\}_{i=1...n}$ ($n$ denotes the task head number). These heads connect with MLLM using the embeddings $\{\mathbf{e}_{i}\}_{i=1...n}$ ($\mathbf{e}_{i}=G(\mathbf{v}_{i})\in\mathbb{R}^{1\times\text{C}}$) transformed from the learnable task-specific tokens $\{\mathbf{v}_{i}\}_{i=1...n}$ ($\mathbf{v}_{i}\in\mathbb{R}^{1\times\text{C}}$) via MLLM.
TPO employs a local-to-global training scheme, first adapting task heads to the MLLM and then training them jointly. Specifically, the MLLM $M$ starts to recognize $\{\mathbf{v}_{i}\}_{i=1...n}$ by updating $G$ according to user instructions, then we tune $\{\mathbf{v}_{i}\}_{i=1...n}$ and $\{H_{i}\}_{i=1...n}$ for adapting visual heads to $M$. Finally, we train both $M$ and $\{H_{i}\}_{i=1...n}$ together. During inference, MLLM-TPO can respond to users’ queries in text, and produce structured visual outputs (like masks, timestamps, trajectories, etc.) when users ask (e.g. yielding time results for “find when the birthday party starts”). We detail how to form MLLM-TPO structurally and train it in the following.
@@ 3.1 Task Preference Models
The task preference model (TPM) $P$ contains a series of task tokens $\{\mathbf{v}_{i}\}_{i=1...n}$ and heads $\{H_{i}\}_{i=1...n}$. Before TPM works, the attached MLLM generates special tokens indicating task types from the input queries. Then TPM dynamically calls the task token $\mathbf{v}_{j}$ based on the special token, transforms it to the task embedding $\mathbf{e}_{j}$ via the LLM $G$ (the last hidden embedding), and feeds $\mathbf{e}_{j}$ to the corresponding visual task head $\{H_{j}\}$ for specific task predictions. Considering existing MLLMs demonstrate remarkable capabilities in common object and scene recognition, yet struggle to accurately locate things or actions, the employed task heads mainly focus on spatiotemporal localization and tracking. Specifically, we give three fundamental task types for compensating mainstream MLLMs’ gaps from expert models in visual perceptions: 1) region head, 2) temporal head, and 3) mask head. Their architectures are given below.
  • Region Head. A two-layer multilayer perceptron (MLP) with ReLU activations is employed for the region head. It takes embeddings from LLM and regresses them to the bounding box coordinates for spatial grounding.
  • Temporal Head. This head is composed of a video encoder, text encoder, and temporal grounding model for moment retrieval and highlight detection. The video and text queries are input into the video and text encoders to obtain their features respectively. Then we append the temporal task embedding after the text features. With the new text and video features, the MLLM estimates the start/end time and highlight score of the corresponding moment from the query via the temporal grounding model.
  • Mask Head. Pixel-level tasks pose significant challenges for MLLMs due to the lack of the corresponding learning data and optimization in MLLMs. To this end, we introduce a specialized mask head, utilizing the image encoder, mask decoder, and memory bank components from SAM2  [68], replacing the prompt encoder with a single MLP layer called the mask adapter. Specifically, for the given visuals and user queries, the features extracted from the aforementioned image encoder and the text representation corresponding to the mask task embeddings from LLM are employed, fed into the mask decoder to produce the desired mask.
Most known discriminative vision tasks can be addressed by one or a combination of these three task heads. TPM builds the architectural foundations for leveraging existing vision data with annotations to enhance MLLMs.
@@ 3.2 Task Preference Optimization
TPO improves MLLMs with extra supervisions from visual task heads by back-propagating gradients from heads to update MLLMs using visual task data. It enables the language model $G$ in MLLM to discriminate specific task types when users demand (task assignment). Then, TPO trains TPM $P$ via compact task representations $\mathbf{e}_{i}$ (task optimization) from $\mathbf{v}_{i}$. Lastly, we train $M$ and $P$ together, tuning $M$ for the refined spatiotemporal perception according to task preferences from $P$. Its optimization objective is given as:
@@
where $\mathbf{T}_{\text{q}}$ and $\mathbf{s}$ denote a user query (some contain specific task needs like tracking) in the form of text token sequence and the ground truth task indicating token, respectively. $\mathbf{A}_{i}$ is the task annotation of the input visual $\mathbf{X}$, so it could be a number tuple for describing regional localization and area or mask for delineating object shape and position. Here $(\mathbf{X},\mathbf{T}_{\text{q}},\mathbf{T}_{\text{a}},\mathbf{A}_{i})$ stands for an input data tuple to MLLM-TPO. Usually, we have $(\mathbf{X},\mathbf{T}_{\text{q}},\mathbf{T}_{\text{a}})$ for typical MLLM training while $(\mathbf{X},\mathbf{T}_{\text{q}},\mathbf{A}_{i})$ for task token and head training. $\mathcal{L}_{\text{assign}}$ is the cross-entropy loss. Meanwhile $\mathcal{L}_{\text{task}_{i}}$ varies according to the task, and is usually regression- or classification-related loss.
To train MLLM-TPO, we propose a 3-stage local-to-global training scheme, given in Figure 3. Stage 1 learns to identify the task type based on user queries. In stage 2, we train task heads along with their corresponding task tokens, respectively. Lastly, we co-train task heads with MLLM by both task data and multimodal conversation data. Our 3-stage training strategy mitigates the risk of degrading the MLLM’s general abilities. We describe them as:
@@
We integrate task tokens and task-specific heads into the model. By training on task-specific data, the model equips the capacity with fine-grained perception and aligns between the task head and the MLLM. The region head, temporal head, mask adapter, and corresponding task tokens are trained respectively. Similarly, the LLM is updated with LoRA.